# XtraLight-MedMamba for Classification of Neoplastic Tubular Adenomas

### v3

@@ 3 Methodology
Figure 2: Architectural structure of XtraLight-MedMamba model for image classification task. The proposed architecture consists of ConvNeXt blocks for local morphological feature extraction with PVM layers for parallel state-space modeling, spatial and channel attention as SCAB modules and FNOClassifier to enforce fixed non-negative orthogonal decision boundaries.
The XtraLight-MedMamba model adopts an architectural framework akin to convolutional neural networks (CNNs), but with a key distinction: instead of relying on convolutional blocks as its primary feature extractors, it employs Parallel Vision Mamba (PVM) layers. The overall architecture comprises six layers, with the number of channels configured as $[8,16,24,32,48,64]$ as shown in Fig. 2. The initial three layers utilize ConvNeXt blocks to extract shallow yet semantically rich features from whole slide image (WSI) tiles. Each ConvNeXt block follows a modernized convolutional design that replaces traditional residual bottlenecks with depthwise separable convolutions, Layer Normalization (LN), and a GELU activation–based MLP-like expansion with a 4:1 channel ratio. These design choices—originally inspired by Transformer architectures—enhance feature expressivity and optimization stability while preserving the strong spatial inductive bias characteristic of convolutional networks. The deeper layers (layers 4 through 6) incorporate PVM layers to capture more complex and nuanced features, by acting as state-space sequence mixers to efficiently model long-range dependencies across spatial regions. This hierarchical combination enables the ConvNeXt front-end to capture localized morphological patterns, while the PVM back-end contextualizes them, globally—bridging low-level texture information with high-level structural understanding critical for differentiating neoplastic from non-neoplastic tissue patterns.
The extracted features from three ConvNext blocks and two PVM layers are fed into a SCAB (spatial and channel attention bridge) module. Adaptive average pooling is performed on feature maps acquired from SCAB and the final PVM layer (stage 6) to standardize spatial dimensions of the feature maps. The pooled features are then concatenated to accumulate and retain the relevant information. The classification head first flattens the high-dimensional feature maps into a one-dimensional vector, effectively aggregating the spatially distributed information extracted by the preceding layers. This flattened vector is then passed through fully connected (dense) layers, which serve to map the learned feature representations to the final output space. In the case of colorectal adenoma classification, this typically corresponds to a set of class probabilities, enabling the model to assign a likelihood score to each potential category for the image: control and cancer.
@@ 3.1 Components of XtraLight-MedMamba Model
3.1.1 ConvNext–Shallow Feature Extractor for XtraLight Model
@@
The ConvNext block [32, 12] builds upon the ResNet-style design by integrating modern architectural refinements inspired by Vision Transformer [3]. Each block begins with a $7\times 7$ depthwise convolution that captures the long-range spatial context within each channel while maintaining computational efficiency through channel grouping, as shown in Fig. 3. The resulting feature map is then permuted from (B, C, H, W) to (B, H, W, C) to enable Layer Normalization and fully connected (linear) transformations that operate along the channel dimension, mimicking the feed-forward network structure of Transformers. This “MLP” sublayer consists of two linear projections separated by a GELU activation, expanding the channel dimension by a factor of mlp_ratio before projecting it back to the original dimensionality. A learnable per-channel scaling parameter $\gamma$ is applied to modulate the transformed output, followed by an inverse permutation that restores the tensor to (B, C, H, W). Finally, a drop path is employed as a form of regularization before adding the residual connection, which preserves gradient flow and stabilizes training. This combination of depth-wise convolution, LayerNorm-based channel mixing, and residual learning enables ConvNeXt blocks to achieve high representational capacity and training stability with minimal computational overhead.
@@ 3.1.2 Parallel Vision Mamba Module
The PVM module [26, 29], as shown on the left side of Fig. 4 (part a), mainly consists of Mamba integrated with residual connections, which improve Mamba’s ability to capture deep spatial relationships. Initially, the input $Y_{in}^{C}$ with channel $C$ goes through layer normalization. The features that are fed into Mamba are first branched out as $X_{1}^{C/4}$, $X_{2}^{C/4}$, $X_{3}^{C/4}$ and $X_{4}^{C/4}$, each with channel $C/4$. The Mamba outputs combined with the residual connection from the inputs and adjustment factor for optimization are then concatenated together to obtain the four feature maps. The right side of Fig. 4 (part b) shows the Mamba component used in the PVM layer. These features are concatenated to obtain $Y_{out}$ with the number of channels as $C$. The concatenated feature outputs are layer normalized and then projected using a Projection operation. By processing deep features in a parallel manner, the PVM module is able to capture multi-scale and intricate features using different kernel sizes. This process significantly reduces the parameters by retaining the same receptive field, which addresses the explosive influence of channel numbers since the number of parameters in Mamba is highly dependent on the number of input channels [29]. The specific operations are expressed by the following equations:
@@
Here, $LN$ denotes the layer normalization, $S_{p}$ is the split operation that splits the input into four separate branches, $\theta$ is the residual scaling factor, catcat denotes the concatenation operation where the outputs are concatenated and $Proj$ is the projection operation where it projects concatenated representation back into the feature space. From Eq. 2, we employed parallel Vision Mamba feature processing while keeping the total number of channels constant, allowing us to preserve high accuracy while achieving significant parameter reduction.
@@ 3.1.3 SCAB Module
The model performance is further improved by the addition of the SCAB module, also known as the Spatial and Channel Attention Bridge [19, 29], for feature propagation. The spatial attention bridge consists of max-pooling, average pooling, and extended convolution of shared weights. The channel attention bridge includes fully connected layers (FCL), global average pooling (GAP), concatenation, and a sigmoid activation function. The SCAB module enhances the sensitivity of the model, ability of the model to converge, and fusion of multi-scale features from different scales [29, 26].
@@ 3.1.4 Fixed Non-Negative Orthogonal Classifier for Model Parameter Reduction
The Fixed Non-Negative Orthogonal Classifier or FNOClassifier [8] module is a lightweight neural classifier designed to map feature embeddings into class logits through structured and normalized projections. During initialization, the model partitions the input feature dimensions into approximately equal groups corresponding to the number of output classes. Random permutation ensures that each class is assigned a unique subset of features, with any remaining dimensions distributed randomly among the classes to maintain balance. A binary base matrix is then constructed, where each row indicates the feature subset associated with a class, and is subsequently L2-normalized along the row dimension to enforce orthogonality and numerical stability. The model also includes an optional input feature normalization step where input vectors are scaled by their L2 norm, ensuring consistent magnitude across samples. In the forward pass, the input which is either normalized or raw is linearly projected using the scaled base matrix, modulated by a learnable scalar parameter $W$ to produce class scores. This allows the FNOClassifier to serve as an efficient and interpretable projection-based classifier, emphasizing structured feature grouping and normalization for robust classification.
Fixed classifiers that incorporate orthogonality are recognized for their cost efficiency and their ability to sometimes outperform learnable classifiers on popular benchmarks. However, existing fixed orthogonal classifiers suffer from geometric limitations that fundamentally prevent them from invoking neural collapse [8, 30].
@@
The FNOClassifier was proposed to resolve the issue that previous fixed orthogonal classifiers fail to invoke neural collapse (NC). Neural collapse is a critical phenomenon in which last-layer features converge to the simplex ETF (Equiangular Tight Frame) structure necessary to achieve global optimality in a layer-peeled model, due to their inherent geometric limitations [8] [30]. The development of the FNOClassifier relies on an analysis of zero-mean NC that specifically considers orthogonality within non-negative Euclidean space. By satisfying the necessary properties of zero-mean NC, the FNOClassifier is designed to induce the optimal solution and maximize the margin of an orthogonal layer-peeled model. This classifier yields an inherent feature dimension separation effect by mitigating feature interference in continual learning and tackling the limitations of mixup on the hypersphere in imbalanced learning, ultimately demonstrating significant performance improvements across various experiments.
@@
where $I_{D}$ is identity matrix and $W_{j,k}$ is the $(j,k)$-th element of $W$. The orthogonality constraint ensures a distinct and independent direction for each class in the feature space, which maximizes the separation between classes and minimizes classifier-induced bias. The non-negativity constraint limits the directions of the classifier to the positive orthant promoting stable and consistent alignment between trained feature representations and their corresponding class-vectors.
@@
where $z(x)$ is the logit vector, $\gamma$ is a positive scaling parameter controlling the magnitude of the logits and $x$ is the input feature representation. $p(d|x)$ is the predicted probability of an input sample with feature representation where $x$ belongs to class $d$.
Neural collapse. Neural collapse refers to a set of geometric properties that emerge toward the end phase of training neural networks using crossentropy loss. First, intra-class collapse occurs. Under zero-mean neural collapse, features from the same class converge toward a shared class mean where $x_{l}^{(d)}$ is the feature embedding of the $l$-th sample belonging to the class $d$, $n_{d}$ and $\mu_{d}$ are the number of samples and the class mean feature vector for class $d$, respectively and defined as:
@@
The convergence over the course of training is denoted by $\xrightarrow[]{\text{training}}$. The condition in (9) indicates within-class feature variance diminishes and the same class feature embeddings concentrate around a single class.
Second, class mean centering occurs. Under zero-mean neural collapse, the weighted mean of the class becomes centered and converges to zero or origin [16], [7], [2]. Therefore, the features behave as:
@@
In (10), $D$ is the total number of classes, the class probability is denoted by $\psi_{d}$ where the global feature distribution is centered at origin and the weighted average of all class means is zero, thus, ensuring a balanced and symmetric arrangement of class means in the feature space.
@@
where all class means have the same Euclidean norm $r$. This constraint ensures equal margins across classes in the feature space by preventing dominance through the larger feature magnitudes. Integrated with intra-class collapse and centered class means, the resulting highly structured feature geometry concentrate samples from each class tightly around their respective means where the class means are evenly distributed around the origin and the angles between class directions are maximized. This yields compact, balanced, and maximally separated class representations consistent with Equiangular Tight Frame (ETF) geometry [16], [33].
@@
Class-Mean Optimization. Under the neural collapse theory, feature embeddings are represented at the level of class means rather than individual samples. Optimization of class mean representations $\mu_{d}$ becomes equivalent to optimization of individual feature vectors as intra-class variance diminishes. Consequently, training with a fixed classifier $W$ amounts to minimizing the cross-entropy over class means $\mu_{d}$:
$\min_{\{\mu_{d}\}}\ \ \mathbb{E}_{d\sim\psi}\!\left[-\log\frac{\exp\big(\gamma\,w_{d}^{\top}\mu_{d}\big)}{\sum_{m=1}^{D}\exp\big(\gamma\,w_{m}^{\top}\mu_{d}\big)}\right]\quad\text{s.t.}\quad\|\mu_{d}\|=r\quad\ \$(12)
@@
where, $d$ is the class index treated as a random variable obtained from the class prior $\psi$ and $\mathbb{E}_{d\sim\psi}[\cdot]$ denotes the average over $d$ under the class distribution $\psi$. $w_{m}$ is the weight vector of the competing class $m$ and $w_{d}$ is the $d$-th column of $W$ for the target class, where $m\neq d$. Under this formulation, optimization solely acts on the feature extractor, aligning each class mean $\mu_{d}$ in the direction of its corresponding classifier while maintaining the separation from the directions associated with other classes. When integrated with equal-norm constraint in (11), it promotes a balanced and symmetric feature geometry consistent with ETF structure of neural collapse [16].
@@
Margin Classification. The objective in 12 is closely related to an implicit max-margin perspective. In certain overparameterized models, training with cross-entropy loss implicitly maximizes the classification margin between classes [23]. Therefore, the classification margin [23] for class $d$ versus class $m$ is expressed in the following margin form:
$\max_{\{\mu_{d}\}}\ \ [\min_{d\neq m}\ \big(w_{d}^{\top}\mu_{d}-w_{m}^{\top}\mu_{d}\big)]\quad\text{s.t.}\quad\|\mu_{d}\|=r\quad$ (13)
@@
where the inner minimization ensures separation from the closest competing class. This drives each class mean $\mu_{d}$ to align with its associated classifier direction $w_{d}$ while maintaining the distance from other class directions. The equal-norm constraint ensures min-margin, max-margin decision boundaries consistent with neural collapse.
@@
Relation to simplex-ETF. In standard learnable classifiers, neural collapse theory predicts that classifier weights align with class means, i.e., $w_{d}\propto\mu_{d}$. The centered and normalized class means form a simplex equiangular tight frame (ETF) $\langle\tilde{\mu}_{d},\tilde{\mu}_{m}\rangle$ $\forall$ $m,d\in\{1,...,D\}$ [33], [16]:

### v4

@@ 3 Methodology
Figure 2: Architectural structure of XtraLight-MedMamba model for image classification task. The proposed architecture comprises ConvNeXt blocks for local morphological feature extraction, PVM layers for parallel state-space modeling, spatial and channel attention as SCAB modules, and an FNOClassifier to enforce fixed, nonnegative, orthogonal decision boundaries.
XtraLight-MedMamba adopts a convolutional neural network (CNN)-like backbone framework, replacing traditional convolution-only feature extractors with ConvNeXt blocks in the early stages and Parallel Vision Mamba (PVM) layers in the later stages as its primary feature extractors. The overall architecture is organized into six sequential stages, with the number of channels set as $[8,16,24,32,48,64]$ as shown in Fig. 2 [39]. The first three stages employ ConvNeXt blocks [46] to generate shallow-to-mid level representations that capture semantically rich features from WSI tiles. Each ConvNeXt block adopts a modernized convolutional design that replaces traditional residual bottlenecks with depthwise separable convolutions, Layer Normalization (LN), and a GELU-activation–based MLP-like expansion with a 4:1 channel ratio. These design choices, inspired by Transformer architectures, enhance feature expressivity and optimization stability while preserving the strong spatial inductive bias characteristic of convolutional networks. The deeper layers, stages four to six, introduce PVM layers [47], [39] that act as state-space sequence mixers, enabling the modeling of long-range dependencies and information integration across spatial regions without the quadratic scaling of self-attention. This hierarchical combination enables the ConvNeXt front-end to capture localized morphological patterns, while the PVM back-end contextualizes them globally, bridging low-level texture information with high-level structural understanding, which is critical for differentiating neoplastic from non-neoplastic tissue patterns.
The multi-level features extracted from the three ConvNeXt blocks, together with the intermediate features from the PVM layers, are fed into a SCAB (spatial and channel attention bridge) module. To standardize the spatial dimensions for classification, feature maps from SCAB and the final PVM layer (stage 6) are passed through adaptive average pooling to achieve a common spatial resolution. The pooled feature maps are then concatenated to form a unified representation that retains relevant information. The resulting high-dimensional feature representation is subsequently flattened into a one-dimensional vector and fed into the FNOClassifier, which produces logits that map the learned embeddings into the final output space. In this study, the classifier outputs a two-class probability distribution corresponding to control and case.
@@ 3.1 Components of XtraLight-MedMamba Model
3.1.1 ConvNeXt–Shallow Feature Extractor for XtraLight-MedMamba Model
@@
The ConvNeXt block [50, 22] builds upon the ResNet-style design by integrating modern architectural refinements inspired by Vision Transformer [8]. Each block begins with a $7\times 7$ depthwise convolution that captures the long-range spatial context within each channel while maintaining computational efficiency through channel grouping, as shown in Fig. 3. The resulting feature map is then permuted from (B, C, H, W) to (B, H, W, C) to enable Layer Normalization and fully connected (linear) transformations that operate along the channel dimension, mimicking the feed-forward network structure of Transformers. This “MLP” sublayer consists of two linear projections separated by a GELU activation, expanding the channel dimension by a factor of mlp_ratio before projecting it back to the original dimensionality. A learnable per-channel scaling parameter $\gamma$ is applied to modulate the transformed output, followed by an inverse permutation that restores the tensor to (B, C, H, W). Finally, a drop_path is employed as a form of regularization before adding the residual connection, thereby preserving gradient flow and stabilizing training. This combination of depth-wise convolution, LayerNorm-based channel mixing, and residual learning enables ConvNeXt blocks to achieve high representational capacity and training stability with minimal computational overhead.
@@ 3.1.2 Parallel Vision Mamba Module
The PVM module [40, 47], as shown on the left side of Fig. 4 (part a) [39], mainly consists of Mamba integrated with residual connections, which improve Mamba’s ability to capture deep spatial relationships. Initially, the input $Y_{in}^{C}$ with channel $C$ goes through layer normalization. The features that are fed into Mamba are first branched out as $X_{1}^{C/4}$, $X_{2}^{C/4}$, $X_{3}^{C/4}$, and $X_{4}^{C/4}$, each with channel $C/4$. The Mamba outputs, combined with the residual connection from the inputs and the optimization adjustment factor, are concatenated to obtain the four feature maps. The right side of Fig. 4 (part b) [39] shows the Mamba component used in the PVM layer. These features are concatenated to obtain $Y_{out}$ with the number of channels as $C$. The concatenated feature outputs are layer-normalized and then projected using a Projection operation. By splitting and processing deep features in parallel, the PVM module can capture multi-scale, intricate features at different granularities before fusion. This process significantly reduces the number of parameters by retaining the same receptive field, thereby avoiding increasing the channel dimension, which is a key consideration given Mamba’s dependence on input dimensionality [47]. The specific operations are expressed by the following equations:
@@
Here, $LN$ denotes the layer normalization, $S_{p}$ is the split operation that splits the input into four separate branches, $\theta$ is the residual scaling factor, catcat denotes the concatenation operation where the outputs are concatenated and $Proj$ is the projection operation where it projects concatenated representation back into the feature space. From Eq. 2, we employed parallel Vision Mamba feature processing while keeping the total number of channels constant, thereby preserving high accuracy while achieving significant parameter reduction.
@@ 3.1.3 SCAB Module
Spatial and Channel Attention Bridge (SCAB) module is incorporated to refine intermediate representations and improve cross-scale feature integration and feature propagation [33, 47]. SCAB module follows two paths: a spatial attention pathway that aggregates max-pooling, average pooling, concatenation operation, followed by extended convolution of shared weights, and a channel attention bridge pathway that includes global average pooling (GAP), concatenation operation, followed by fully connected layers (FCL), and sigmoid activation function. Together, this attention-based fusion improves the model’s sensitivity, optimization stability, and multi-scale feature propagation [47, 40].
@@ 3.1.4 Fixed Non-Negative Orthogonal Classifier for Model Parameter Reduction
The Fixed Non-Negative Orthogonal Classifier (FNOClassifier) [16] module is a lightweight neural classifier that maps feature embeddings to class logits via structured, normalized projections. During initialization, the model partitions the input feature dimensions into approximately equal groups, each corresponding to an output class. A random permutation ensures that each class is assigned a unique subset of features, with any remaining dimensions randomly distributed among the classes to maintain balance. A binary base matrix is then constructed, where each row indicates the feature subset associated with a class, and is subsequently L2-normalized along the row dimension to enforce orthogonality and numerical stability. The model also includes an optional input feature normalization step, in which input vectors are scaled by their L2 norms, ensuring consistent magnitude across samples. In the forward pass, the input, which is either normalized or raw, is linearly projected using the scaled base matrix, modulated by a learnable scalar parameter $W$ to produce class scores. This allows the FNOClassifier to serve as an efficient and interpretable projection-based classifier, emphasizing structured feature grouping and normalization for robust classification.
Fixed classifiers that incorporate orthogonality are recognized for their cost efficiency and, in some cases, their ability to outperform learnable classifiers on popular benchmarks. However, existing fixed orthogonal classifiers suffer from geometric limitations that fundamentally prevent them from invoking neural collapse [16, 48].
@@
The FNOClassifier was proposed to resolve the issue that previous fixed orthogonal classifiers fail to invoke neural collapse (NC). NC is a critical phenomenon in which last-layer features converge to the simplex ETF (Equiangular Tight Frame) structure necessary to achieve global optimality in a layer-peeled model, due to their inherent geometric limitations [16] [48]. The development of the FNOClassifier relies on an analysis of zero-mean NC that explicitly accounts for orthogonality in non-negative Euclidean space. By satisfying the necessary properties of zero-mean NC, the FNOClassifier is designed to induce an optimal solution and maximize the margin of an orthogonal-layer peeled model. This classifier yields an inherent feature-dimension separation effect by mitigating feature interference in continual learning and addressing the limitations of mixup on the hypersphere in imbalanced learning, ultimately demonstrating significant performance improvements across various experiments.
@@
where $I_{D}$ is identity matrix, $W_{j,k}$ is the $(j,k)$-th element of $W$ and $W^{\top}$ is the transpose of $W$. The orthogonality constraint ensures that each class has a distinct, independent direction in feature space, thereby maximizing inter-class separation and minimizing classifier-induced bias. The non-negativity constraint restricts the classifier to the positive orthant, promoting stable, consistent alignment between trained feature representations and their corresponding class vectors.
@@
where $z(x)$ is the logit vector, $\gamma$ is a positive scaling parameter controlling the magnitude of the logits, and $x$ is the input feature representation. $p(d|x)$ is the predicted probability of an input sample with feature representation where $x$ belongs to class $d$.
Neural collapse. Neural collapse refers to a set of geometric properties that emerge toward the end phase of training neural networks using cross-entropy loss. First, intra-class collapse occurs. Under zero-mean neural collapse, features from the same class converge toward a shared class mean, where $x_{l}^{(d)}$ is the feature embedding of the $l$-th sample belonging to the class $d$, $n_{d}$ and $\mu_{d}$ are the number of samples and the class mean feature vector for class $d$, respectively and defined as:
@@
The convergence over the course of training is denoted by $\xrightarrow[]{\text{training}}$. The condition in (9) indicates that within-class feature variance diminishes, and the same class feature embeddings concentrate around a single class.
Second, class means centering occurs. Under zero-mean neural collapse, the weighted mean of the class becomes centered and converges to zero or the origin [27], [14], [6]. Therefore, the features behave as:
@@
In (10), $D$ is the total number of classes, the class probability is denoted by $\psi_{d}$, where the global feature distribution is centered at the origin and the weighted average of all class means is zero, thus, ensuring a balanced and symmetric arrangement of class means in the feature space.
@@
where all class means have the same Euclidean norm $r$. This constraint ensures equal margins across classes in feature space by preventing dominance due to larger feature magnitudes. Integrated with intra-class collapse and centered class means, the resulting highly structured feature geometry concentrates samples from each class tightly around their respective means, where the class means are evenly distributed around the origin, and the angles between class directions are maximized. This yields compact, balanced, and maximally separated class representations consistent with Equiangular Tight Frame (ETF) geometry [27], [51].
@@
Class-Mean Optimization. Under the NC theory, feature embeddings are represented at the level of class means rather than individual samples. Optimization of class mean representations $\mu_{d}$ becomes equivalent to optimization of individual feature vectors as intra-class variance diminishes. Consequently, training with a fixed classifier $W$ amounts to minimizing the cross-entropy over class means $\mu_{d}$:
$\min_{\{\mu_{d}\}}\left[\mathbb{E}_{d\sim\psi}\!\left[-\log\frac{\exp\big(\gamma\,w_{d}^{\top}\mu_{d}\big)}{\sum_{m=1}^{D}\exp\big(\gamma\,w_{m}^{\top}\mu_{d}\big)}\right]\right]\quad\text{s.t.}\quad\|\mu_{d}\|=r\quad\ \$(12)
@@
where $d$ is the class index treated as a random variable obtained from the class prior $\psi$ and $\mathbb{E}_{d\sim\psi}[\cdot]$ denotes the average over $d$ under the class distribution $\psi$. $w_{m}$ is the weight vector of the competing class $m$ and $w_{d}$ is the $d$-th column of $W$ for the target class, where $m\neq d$. Under this formulation, optimization acts only on the feature extractor, aligning each class mean $\mu_{d}$ with its corresponding classifier while maintaining separation from the directions associated with other classes. When integrated with an equal-norm constraint in (11), it promotes a balanced and symmetric feature geometry consistent with the ETF structure of neural collapse [27].
@@
Margin Classification. The objective in 12 is closely related to an implicit max-margin perspective. In certain over-parameterized models, training with cross-entropy loss implicitly maximizes the classification margin between classes [38]. Therefore, the classification margin [38] for class $d$ versus class $m$ is expressed in the following margin form:
$\max_{\{\mu_{d}\}}\left[\min_{d\neq m}\big(w_{d}^{\top}\mu_{d}-w_{m}^{\top}\mu_{d}\big)\right]\quad\text{s.t.}\quad\|\mu_{d}\|=r\quad$ (13)
@@
where the inner minimization ensures separation from the closest competing class. This drives each class means $\mu_{d}$ to align with its associated classifier direction $w_{d}$ while maintaining a distance from the other class directions. The equal-norm constraint ensures that min-margin and max-margin decision boundaries are consistent with NC.
@@
Relation to simplex-ETF. In standard learnable classifiers, NC theory predicts that classifier weights align with class means, i.e., $w_{d}\propto\mu_{d}$. The centered and normalized class means form a simplex equiangular tight frame (ETF) $\langle\tilde{\mu}_{d},\tilde{\mu}_{m}\rangle$ $\forall$ $m,d\in\{1,...,D\}$ [51], [27]: